OpenStudy (anonymous):

what is the standard form equation of a line passes through (5,-3) with a slope of -3

OpenStudy (anonymous):

y=mx+b where m is the slope and b is the y intercept

OpenStudy (anonymous):

thus u can use y-y1 = m(x-x1) to get the line so... y - (-3) = -3(x - 5)

OpenStudy (anonymous):

A/ 3x+y=12 B/3x+y=-12 C/3x-y=12 D/3x-y=-12

OpenStudy (anonymous):

ahh lol

OpenStudy (anonymous):

probably A the line of the equation is y = -3x+12

OpenStudy (anonymous):

thus 3x+y = 12

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

:)